OpenStudy (anonymous):

Which of the following is not a perfect square trinomial? A. 81 + 18y + y^2 B. 100 – 20y + y^2 C. 144 + 24y + y^2 D. 121 + 11y + y^2

OpenStudy (anonymous):

@agent0smith @mathstudent55 @LadyKatsaTheGraceling @Nurali

OpenStudy (the_fizicx99):

d

OpenStudy (nurali):

i think 121 +11y +y^2 = y^2 +11y +121 = *** does not factor ****

OpenStudy (the_fizicx99):

It comes in the form of (a-b)^2 = a^2 - 2(a)(b) + b^2

OpenStudy (anonymous):

thanks :) @Nurali @tHe_FiZiCx99

OpenStudy (the_fizicx99):

yw

OpenStudy (nurali):

Anytime.

OpenStudy (anonymous):

I have another one kind of similar to the other one if you can help me with that??

OpenStudy (the_fizicx99):

Sure

OpenStudy (anonymous):

okay Factor. m^2 + 12m + 36 A. Cannot be factored B. (m + 36)^2 C. (m – 6)^2 D. (m + 6)^2

OpenStudy (the_fizicx99):

Going with D

OpenStudy (anonymous):

wow cool thanks :) I might have one more?

OpenStudy (the_fizicx99):

Have to get off :/ gl

OpenStudy (anonymous):

haha it's ok I didn't end up having another one anyways, but thanks!! :)